# matrix,arrows,decorations.pathmorphing

### v2

@@ Abstract.
We show that every automorphism of the congruence completion of the extended mapping class group which preserves the set of conjugacy classes of procyclic groups generated by Dehn twists is inner and that its automorphism group is naturally isomorphic to the automorphism group of the procongruence pants complex. In the genus 00 case, we prove the stronger result that all automorphisms of the profinite completion of the extended mapping class group are inner.
AMS Math Classification: 14G32, 11R32, 14D23, 20F34, 57M10.
@@ 1. Introduction
Let $S=S_{g,n}$ be a orientable surface of genus $g(S)=g$ with $n(S)=n$ punctures. We will assume that $S$ has negative Euler characteristic: $\chi(S)=2-2g-n<0$. Let then ${\Gamma}^{\pm}(S)$ be the extended mapping class group of $S$, namely the group of isotopy classes of diffeomorphisms of $S$ and ${\Gamma}(S)$ be the mapping class group of $S$, i.e. the subgroup of ${\Gamma}^{\pm}(S)$ consisting of the mapping classes which preserve the orientation of $S$. We denote, respectively, by $\mathrm{P}\Gamma(S)$ and $\mathrm{P}\Gamma^{\pm}(S)$ the pure mapping class group and the pure extended mapping class group, namely the subgroups of ${\Gamma}(S)$ and ${\Gamma}^{\pm}(S)$ consisting of those mapping classes which fix pointwise the punctures. We also let $d(S)=3g-3+n$ be the modular dimension of $S$, that is to say the dimension of the moduli stack ${\mathcal{M}}(S)$ of smooth curves whose complex models are diffeomorphic to $S$.
@@
The proof of all the above results is based on the study of the complex of curves $C(S)$. This is the abstract simplicial complex of dimension $d(S)-1$ whose simplices are the sets of isotopy classes of essential simple closed curves on $S$ which admit disjoint representatives (such sets are called multicurves). There is a natural action of the extended mapping class ${\Gamma}^{\pm}(S)$ on $C(S)$ and the above results are obtained showing that all automorphisms of $C(S)$ are induced by this action. In fact, this action corresponds to the inner action of ${\Gamma}^{\pm}(S)$ on the set of Dehn twists, which are parametrized by the vertices of $C(S)$, and, if we denote by $\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))$ the group of automorphisms of ${\Gamma}^{\pm}(S)$ which preserve the set of conjugacy classes of Dehn twists, there is a series of inclusion:
@@
which are showed to be all isomorphisms for $d(S)>1$ (with the only exception of the case $g(S)=1$ and $n(S)=2$). The identity $\operatorname{Inn}({\Gamma}^{\pm}(S))=\operatorname{Aut}({\Gamma}^{\pm}(S))$ then follows from the fact that all automorphisms of ${\Gamma}^{\pm}(S)$, for $d(S)>1$, with few low genus exceptions, preserve the set of conjugacy classes of Dehn twists.
@@
In [19], Margalit determined the automorphism group of a related complex, the so called pants graph $C_{P}(S)$. This is defined as follows. The vertices of $C_{P}(S)$ are pants decompositions (i.e. maximal multicurves) of $S$ and correspond to facets (simplices of highest dimension $d(S)-1$) of $C(S)$. Two vertices $\underline{{\alpha}}=(\alpha_{0},\alpha_{1},\ldots,\alpha_{d(S)-1})$ and $\underline{{\alpha}}^{\prime}=(\alpha^{\prime}_{0},\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{d(S)-1})$ are connected by an edge if they differ by an elementary move, that is to say: the multi-curves $\underline{{\alpha}}$ and $\underline{{\alpha}}^{\prime}$ have $d(S)-1$ elements in common, so that, up to relabelling, ${\alpha}_{i}={\alpha}^{\prime}_{i}$, for $i=1,\ldots,d(S)-1$, and the surface $S^{\prime}$ obtained cutting $S$ along the curves ${\alpha}_{i}$, for $i>0$, is a surface of modular dimension $1$, i.e. $S^{\prime}=S_{1,1}$ or $S^{\prime}=S_{0,4}$. Then, ${\alpha}_{0}$ and ${\alpha}^{\prime}_{0}$, which are supported on $S^{\prime}$, should intersect in a minimal way, i.e. they have geometric intersection number $1$, in the first case, and $2$, in the second case.
@@
Margalit’s results (cf. Theorem 1 and 2 in [19]) then imply that, for $d(S)>1$, the natural action of ${\Gamma}^{\pm}(S)$ on $C_{P}(S)$ induces a series of isomorphisms:
@@
(1) $\operatorname{Inn}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}(C_{P}(S)).$
@@
In the paper [6], we proved a partial analogue of the above series of isomorphisms in the setting of procongruence mapping class groups which we now proceed to define.
@@
The profinite mapping class groups ${\widehat{\Gamma}}^{\pm}(S)$, ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}(S)$ are defined to be the profinite completions of ${\Gamma}^{\pm}(S)$, ${\Gamma}(S)$, $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}\Gamma(S)$, respectively. The procongruence mapping class groups ${\check{\Gamma}}^{\pm}(S)$, ${\check{\Gamma}}(S)$, $\mathrm{P}{\check{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\check{\Gamma}}(S)$ are the images of ${\widehat{\Gamma}}^{\pm}(S)$, ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}(S)$, respectively, in the profinite group $\operatorname{Out}(\widehat{\pi_{1}(S)})$, where, for an abstract group $G$, we denote by $\widehat{G}$ its profinite completion. By a classical result of Grossman, the natural homomorphism from each of the above mapping class group to either its profinite or procongruence completion is injective and we then identify the abstract groups with their images in the corresponding profinite groups. By the results of [3] and [5], the profinite and procongruence completion of mapping class groups coincide if $g(S)\leq 2$. For $g(S)\geq 3$, this is an open problem. We will then rather stick with the procongruence completion, since, in contrast with the profinite completion, some basic combinatorial properties are known, thanks to the results contained in [4] and [5].
@@
In analogy with the topological case, for the study of the automorphism group of procongruence mapping class groups, it is useful to introduce the procongruence curve complex ${\check{C}}(S)$. This is an abstract simplicial profinite complex (cf. Definition 3.2 in [4]) of dimension $d(S)-1$, naturally associated to the congruence completion of the mapping class group and endowed with a natural continuous action of ${\check{\Gamma}}(S)$. Roughly speaking (for the precise definition we refer the reader to Section 4 of [4] or Section 4.6 in [6]), ${\check{C}}(S)$ is obtained as the inverse limit of the quotients of $C(S)$ by the action of congruence levels of ${\Gamma}(S)$, that is to say finite index subgroups of ${\Gamma}(S)$ which are open for the congruence topology.
@@
The set of profinite Dehn twists of $\mathrm{P}{\check{\Gamma}}(S)$ is defined to be the closure, inside this group, of the set of Dehn twists of $\mathrm{P}\Gamma(S)\subset\mathrm{P}{\check{\Gamma}}(S)$. The key property of ${\check{C}}(S)$ is then that its set of $k$-simplices parameterizes the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{k}\}$ of closed abelian subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ of rank $k+1$, topologically generated by the profinite Dehn twists of $\mathrm{P}{\check{\Gamma}}(S)$ (cf. Theorem 6.9 in [4]). Here, for ${\sigma}\in{\check{C}}(S)$, we let ${\hat{\mathrm{I}}}_{\sigma}$ be the closed abelian subgroup of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by the profinite Dehn twists parameterized by the vertices of ${\sigma}$.
@@
The natural action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}(S)$ then corresponds to the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on the set of abelian groups parameterized by ${\check{C}}(S)$ and, if we denote by $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ (resp. $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$) the group of those automorphisms of ${\check{\Gamma}}^{\pm}(S)$ (resp. $\mathrm{P}{\check{\Gamma}}(S)$) which preserve the set of conjugacy classes of the procyclic subgroups generated by profinite Dehn twists, we get a series of natural monomorphism:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))\hookrightarrow\operatorname{Aut}({\check{C}}(S)).$
@@
The procongruence pants complex ${\check{C}}_{P}(S)$ is the profinite graph which is, roughly speaking (cf. Section 6.2 in [6] for the precise definition), the inverse limit of the quotients of $C_{P}(S)$ by the action of congruence levels of ${\Gamma}(S)$. It is also endowed, by definition, with a natural continuous action of ${\check{\Gamma}}^{\pm}(S)$. The profinite set of vertices of ${\check{C}}_{P}(S)$ identifies with the profinite set of facets of ${\check{C}}(S)$ and so it parameterizes the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$ of maximal abelian subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by profinite Dehn twists. The natural continuous action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}_{P}(S)$ is then induced by the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on this profinite set.
The main result of the paper is an analogue of the series of isomorphisms (1):
@@
For a connected hyperbolic surface $S$ such that $d(S)>1$, there is a series of natural isomorphisms:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S)).$
@@
The condition on automorphisms in Theorem 1.1 is slightly more restrictive than the one considered in [6], where we denoted by $\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ the subgroup of $\operatorname{Aut}({\check{\Gamma}}(S))$ consisting of those automorphisms which preserve the conjugacy classes of decomposition subgroups of ${\check{\Gamma}}(S)$ (cf. Definition 7.1 in [6]). It is not difficult to see that there is an inclusion $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))\subseteq\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ which for $d(S)>1$ and $Z({\check{\Gamma}}(S))=\{1\}$ is indeed an equality but otherwise is strict.
In [6], we defined the arithmetic procongruence mapping class group ${\check{\Gamma}}_{\mathbb{Q}}(S)$ to be the image of the étale fundamental group of the moduli stack ${\mathcal{M}}(S)_{\mathbb{Q}}:={\mathcal{M}}(S)\times\operatorname{Spec}(\mathbb{Q})$ by the monodromy representation associated to the universal curve over ${\mathcal{M}}(S)_{\mathbb{Q}}$. We then proved that, for $d(S)>1$, there holds (cf. Theorem 9.16 in [6]):
$\operatorname{Inn}({\check{\Gamma}}_{\mathbb{Q}}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}_{\mathbb{Q}}(S)).$
The identity $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ in Theorem 1.1 is the real analogue of this result. The extended procongruence mapping class group ${\check{\Gamma}}^{\pm}(S)$ can indeed be defined to be the image of the étale fundamental group of the moduli stack ${\mathcal{M}}(S)_{\mathbb{R}}:={\mathcal{M}}(S)\times\operatorname{Spec}(\mathbb{R})$ by the monodromy representation associated to the universal curve over ${\mathcal{M}}(S)_{\mathbb{R}}$.
@@
Theorem 1.1 is more striking than Theorem 9.16 in [6], since ${\check{\Gamma}}^{\pm}(S)$ contains ${\check{\Gamma}}(S)$ as a normal index $2$ subgroup, rather than as a normal subgroup of infinite index, and we know that $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ is quite far from consisting only of inner automorphisms. In fact, the outer automorphism group $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$ contains a copy of the absolute Galois group of the rationals $G_{\mathbb{Q}}$ (cf. Corollary 7.6 in [4]).
@@
The analogy between Theorem 1.1 and Margalit’s series of isomorphisms (1) falls short only in that, in contrast with the topological case, we do not know whether, for $d(S)>1$ and $Z({\check{\Gamma}}(S))=\{1\}$, there holds $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}({\check{\Gamma}}^{\pm}(S))$. However, thanks to a recent result by Hoshi, Minamide and Mochizuki (cf. [11]), we are able to fill this gap in genus 00. Since, in this case, we also know that the congruence subgroup property holds, we get:

### v3

# Automorphisms of the procongruence pants complex

@@ Abstract.
We show that every automorphism of the congruence completion of the extended mapping class group that preserves the set of conjugacy classes of procyclic groups generated by Dehn twists is inner, and that its automorphism group is naturally isomorphic to the automorphism group of the procongruence pants complex. In the genus-zero case, we prove the stronger result that all automorphisms of the profinite completion of the extended mapping class group are inner.
AMS Math Classification: 14G32, 11R32, 20F34, 14D23, 57M10.
@@ 1. Introduction
Let $S=S_{g,n}$ be a closed orientable surface of genus $g(S)=g$ with $n(S)=n$ punctures. We will assume that $S$ has negative Euler characteristic: $\chi(S)=2-2g-n<0$. Let ${\Gamma}^{\pm}(S)$ be the extended mapping class group of $S$, namely the group of isotopy classes of diffeomorphisms of $S$, and ${\Gamma}(S)$ be the mapping class group of $S$, i.e. the subgroup of ${\Gamma}^{\pm}(S)$ consisting of the mapping classes which preserve the orientation of $S$. We denote respectively by $\mathrm{P}\Gamma(S)$ and $\mathrm{P}\Gamma^{\pm}(S)$ the pure mapping class group and the pure extended mapping class group, namely the subgroups of ${\Gamma}(S)$ and ${\Gamma}^{\pm}(S)$ consisting of those mapping classes which fix pointwise the punctures.
Let ${\mathcal{M}}(S)$ be the moduli stack of smooth curves whose complex models are diffeomorphic to $S$. This is a smooth Deligne–Mumford (briefly DM) stack over $\operatorname{Spec}\mathbb{Z}$ with the property that the topological fundamental group of ${\mathcal{M}}(S)_{\mathbb{C}}:={\mathcal{M}}(S)\times\operatorname{Spec}\mathbb{C}$ identifies with ${\Gamma}(S)$.
For $S=S_{g,n}$, we will sometimes also denote ${\mathcal{M}}(S)$ by ${\mathcal{M}}_{g,[n]}$. We will instead denote by ${\mathcal{M}}_{g,n}$ the étale covering of ${\mathcal{M}}_{g,[n]}$ obtained fixing an order on the punctures of $S$. The topological fundamental group of $({\mathcal{M}}_{g,n})_{\mathbb{C}}$ then identifies with $\mathrm{P}\Gamma(S_{g,n})$. We let $d(S):=\dim{\mathcal{M}}(S)=3g-3+n$ and call this number the modular dimension of $S$.
@@
The proof of all the above results is based on the study of the complex of curves $C(S)$. This is the abstract simplicial complex of dimension d(S)1d(S)-1 whose simplices are the sets of isotopy classes of essential simple closed curves on $S$ which admit disjoint representatives (such sets are called multicurves). There is a natural action of the extended mapping class group ${\Gamma}^{\pm}(S)$ on $C(S)$ and the above results are obtained showing that all automorphisms of $C(S)$ are induced by this action. In fact, this action corresponds to the inner action of ${\Gamma}^{\pm}(S)$ on the set of Dehn twists, which are parametrized by the vertices of $C(S)$, and, if we denote by $\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))$ the group of automorphisms of ${\Gamma}^{\pm}(S)$ which preserve the set of conjugacy classes of Dehn twists, there is a series of monomorphisms:
@@
which, for $d(S)>1$, with the only exception of the case $g(S)=1$ and $n(S)=2$, are all shown to be isomorphisms. The identity $\operatorname{Inn}({\Gamma}^{\pm}(S))=\operatorname{Aut}({\Gamma}^{\pm}(S))$ then follows from the fact that all automorphisms of ${\Gamma}^{\pm}(S)$, for $d(S)>1$, with few low-genus exceptions, preserve the set of conjugacy classes of Dehn twists.
@@
In [29], Margalit determined the automorphism group of a related complex, the so-called pants graph $C_{P}(S)$. This is defined as follows: the vertices of $C_{P}(S)$ are pants decompositions (i.e. maximal multicurves) of $S$ and correspond to facets (simplices of highest dimension d(S)1d(S)-1) of $C(S)$. Two vertices $\underline{{\alpha}}=(\alpha_{0},\alpha_{1},\ldots,\alpha_{d(S)-1})$ and $\underline{{\alpha}}^{\prime}=(\alpha^{\prime}_{0},\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{d(S)-1})$ are connected by an edge if they differ by an elementary move, that is: the multicurves $\underline{{\alpha}}$ and $\underline{{\alpha}}^{\prime}$ have d(S)1d(S)-1 elements in common, so that, up to relabeling ${\alpha}_{i}={\alpha}^{\prime}_{i}$, for $i=1,\ldots,d(S)-1$, and the surface $S^{\prime}$ obtained cutting $S$ along the curves ${\alpha}_{i}$, for $i>0$, is a surface of modular dimension $1$, i.e. $S^{\prime}=S_{1,1}$ or $S^{\prime}=S_{0,4}$. Then, ${\alpha}_{0}$ and ${\alpha}^{\prime}_{0}$, which are supported on $S^{\prime}$, should intersect in a minimal way, i.e. they have geometric intersection number $1$, in the first case, and $2$ in the second case.
@@
Margalit’s results (cf. [29, Theorem 1 and 2]) then imply that, for $d(S)>1$, the natural action of ${\Gamma}^{\pm}(S)$ on $C_{P}(S)$ induces a series of isomorphisms:
(1) $\operatorname{Inn}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}(S))\cong\operatorname{Aut}(C_{P}(S)).$
When the center $Z({\Gamma}(S))$ is trivial, that is, for $d(S)>1$ and $S\neq S_{1,2}$ and $S_{2}$, there is also a series of isomorphisms:
(2) $\operatorname{Aut}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}(S))\cong\operatorname{Aut}({\Gamma}(S)).$
For $S=S_{1,1},S_{1,2}$ or $S_{2}$, the center $Z({\Gamma}(S))=Z({\Gamma}^{\pm}(S))$ is instead generated by the hyperelliptic involution $\upsilon$} and so $Z({\Gamma}(S))\cong\mathbb{Z}/2$. In this case, there are two epimorphisms $\xi\colon\thinspace{\Gamma}^{\pm}(S)/Z({\Gamma}(S))\to Z({\Gamma}(S))$ and $\zeta\colon\thinspace{\Gamma}^{\pm}(S)/Z({\Gamma}(S))\to Z({\Gamma}(S))$, where the latter is induced by the orientation character (so that its restriction to ${\Gamma}(S)/Z({\Gamma}(S))$ is trivial), which determine the automorphisms $\exp\xi$ and $\exp\zeta$ of ${\Gamma}^{\pm}(S)$, defined by the assignments $x\mapsto x\cdot\xi(\bar{x})$ and $x\mapsto x\cdot\zeta(\bar{x})$, where $\bar{x}$ denotes the image of $x$ in the quotient ${\Gamma}^{\pm}(S)/Z({\Gamma}(S))$. In these cases, we have the isomorphisms (cf. Section 3.6):
@@
(3) $\operatorname{Aut}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))\times\langle\exp\xi\rangle\hskip 14.22636pt\mbox{and}\hskip 14.22636pt\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\Gamma}(S))\times\langle\exp\zeta\rangle.$
@@
In the paper [9], we proved a partial analogue of the series of isomorphisms (1) in the setting of procongruence mapping class groups which we now proceed to define.
The profinite mapping class groups ${\widehat{\Gamma}}^{\pm}(S)$, ${\widehat{\Gamma}}(S)$, $\mathrm{P}{\widehat{\Gamma}}^{\pm}(S)$ and $\mathrm{P}{\widehat{\Gamma}}(S)$ are defined to be the profinite completions of ${\Gamma}^{\pm}(S)$, ${\Gamma}(S)$, $\mathrm{P}\Gamma^{\pm}(S)$ and $\mathrm{P}\Gamma(S)$, respectively.
[…]
@@
The profinite and the procongruence completions of the mapping class group coincide for $g(S)\leq 2$ (cf. [3], [30], [5] and [8]), while, for $g(S)\geq 3$, this is an open problem. For this reason, we will stick with the procongruence completion, since, in contrast with the profinite completion, some basic combinatorial properties are known, thanks to the results contained in [6] and [8].
@@
In analogy with the topological case, for the study of the automorphism group of procongruence mapping class groups, it is useful to introduce the procongruence curve complex ${\check{C}}(S)$. This is a simplicial profinite complex (cf. [6, Definition 3.2]) of dimension d(S)1d(S)-1, naturally associated to the procongruence completion of the mapping class group and endowed with a natural continuous action of ${\check{\Gamma}}(S)$. Naively, ${\check{C}}(S)$ can be described as the inverse limit of the quotients of $C(S)$ by the action of congruence levels of ${\Gamma}(S)$, that is, finite index subgroups of ${\Gamma}(S)$ which are open for the congruence topology.
However, this is not how ${\check{C}}(S)$ is actually defined, and for the precise (rather technical) definition, we urge the reader to look at [6, Section 4] or [9, Section 4.6]. In particular, ${\check{C}}(S)$ is a genuine abstract simplicial complex so that, for instance, all the usual definitions of star and link of a simplex in a simplicial complex make perfect sense for it and do not need an ad hoc definition.
The set of profinite Dehn twists of $\mathrm{P}{\check{\Gamma}}(S)$ is the closure, inside this group, of the set of Dehn twists of $\mathrm{P}\Gamma(S)\subset\mathrm{P}{\check{\Gamma}}(S)$. These elements were first introduced in [4] for profinite mapping class groups and studied extensively in [6] for procongruence mapping class groups. A similar and obviously related notion for étale fundamental groups of configuration spaces of points on algebraic curves has been later introduced in [17].
The key property of ${\check{C}}(S)$ is then that its set of $k$-simplices parameterizes the inertia groups of $\mathrm{P}{\check{\Gamma}}(S)$ (cf. [6, Theorem 6.9]):
Definition 1.1.
@@
For a simplex ${\sigma}\in{\check{C}}(S)$, we define ${\hat{\mathrm{I}}}_{\sigma}$ to be the closed abelian subgroup of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by the profinite Dehn twists parameterized by the vertices of $\sigma$}. This is called the inertia group associated to $\sigma$}.
Note that the natural action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}(S)$ corresponds to the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on the profinite set of inertia groups:
Definition 1.2.
@@
Let $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$ (resp. $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ and $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$) be the subgroup of $\operatorname{Aut}(\mathrm{P}{\check{\Gamma}}(S))$ (resp. $\operatorname{Aut}({\check{\Gamma}}(S))$ and $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$) consisting of those automorphisms which preserve the set of conjugacy classes of procyclic inertia groups.
Remark 1.3.
From [9, Proposition 7.2], it follows that the elements of $\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))$, $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ and $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))$ preserve the set of conjugacy classes of all inertia groups.
There is then a series of natural monomorphisms:
$\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\subseteq\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))\hookrightarrow\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}(S))\hookrightarrow\operatorname{Aut}({\check{C}}(S)).$
@@
The procongruence pants complex ${\check{C}}_{P}(S)$ is the $1$-dimensional simplicial profinite complex which realizes the inverse limit of the quotients of $C_{P}(S)$ by the action of congruence levels of ${\Gamma}(S)$ (cf. [9, Section 6.2], for the precise definition). It is also endowed, by definition, with a natural continuous action of ${\check{\Gamma}}^{\pm}(S)$. The profinite set of vertices of ${\check{C}}_{P}(S)$ identifies with the profinite set of facets of ${\check{C}}(S)$ and so it parameterizes the profinite set $\{{\hat{\mathrm{I}}}_{\sigma}|\,{\sigma}\in{\check{C}}(S)_{d(S)-1}\}$ of maximal abelian subgroups of $\mathrm{P}{\check{\Gamma}}(S)$ topologically generated by profinite Dehn twists. The natural continuous action of ${\check{\Gamma}}^{\pm}(S)$ on ${\check{C}}_{P}(S)$ is then induced by the conjugacy action of ${\check{\Gamma}}^{\pm}(S)$ on this profinite set.
The main result of the paper is an analogue of the series of isomorphisms (1) and (2) and of the identities (3):
@@
For $S$ a connected hyperbolic surface, we have:
  1. (i)
    There is an isomorphism $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S))$.
  2. (ii)
    For $d(S)>1$ and $S\neq S_{1,2},S_{2}$, there is a series of natural isomorphisms:
    $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}^{\mathbb{I}}(\mathrm{P}{\check{\Gamma}}^{\pm}(S)).$
  3. (iii)
    For $S=S_{1,2}$ or $S_{2}$, there holds:
    $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\times\langle\exp\zeta\rangle.$
@@
The condition on automorphisms in Theorem 1.4 is slightly more restrictive than the one considered in [9], where we denoted by $\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ the subgroup of $\operatorname{Aut}({\check{\Gamma}}(S))$ consisting of those automorphisms which preserve the conjugacy classes of decomposition subgroups of ${\check{\Gamma}}(S)$ (cf. [9, Definition 7.1]). It is not difficult to see that there is an inclusion $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))\subseteq\operatorname{Aut}^{\ast}({\check{\Gamma}}(S))$ which for $d(S)>1$ and $Z({\check{\Gamma}}(S))=\{1\}$ is indeed an equality but otherwise is strict.
Throughout this article, we will make use of several results established in our former paper [9], which studies the procongruence curve and pants complexes and has a large overlap with the article [26], for the reasons explained there. As the referee pointed out to us, other notions of Dehn twists and curve complexes appeared in [14, 25, 41], both in the profinite and pro-$p$ context.
Let us also observe, answering another of the referee’s queries, that the isomorphism $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))\cong\operatorname{Aut}({\check{C}}_{P}(S))$ of (i), Theorem 1.4 is entirely different from the (only apparently) similar isomorphism stated in [26, Theorem 7.1]. In fact, in Lochak’s statement the procongruence graph ${\check{C}}_{P}(S)$ is replaced by the pro-object ${\check{C}}_{P}(S)_{st}$ determined by all orbifold quotients of the 2-dimensional pants complex by finite index subgroups of ${\Gamma}(S)$ which are open for the congruence topology. Now, if $X$ denotes the 2-dimensional pants complex and $X^{\lambda}$ one of such orbifold quotients, since $X$ is simply connected, every automorphism of $X^{\lambda}$ lifts to an automorphism of $X$, so that [26, Theorem 7.1] is a formal consequence (cf. [26, Section 7.4]) of Margalit’s rigidity theorem (cf. [29, Theorem 1]).
Note that the first item of Theorem 1.4 is just an improvement, though a substantial one, of [9, Theorem 8.1]. On the other hand, the isomorphisms in the second and third items of Theorem 1.4 are surprising new results and should be regarded as the main achievement of this paper. In fact, in the topological case, for a connected hyperbolic surface $S$ such that $d(S)>1$ and $S\neq S_{1,2},S_{2}$, there is a series of identities:
$\operatorname{Aut}^{\mathbb{I}}({\Gamma}(S))=\operatorname{Inn}({\Gamma}^{\pm}(S))=\operatorname{Aut}^{\mathbb{I}}({\Gamma}^{\pm}(S)).$
@@
However, there is no analogous result for the procongruence mapping class group ${\check{\Gamma}}(S)$, since its automorphism group $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ is quite far from only consisting of the automorphisms induced by restriction of inner automorphisms of ${\check{\Gamma}}^{\pm}(S)$. As a matter of fact, the outer automorphism group $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$ contains a copy of the absolute Galois group of the rationals $G_{\mathbb{Q}}$ (cf. [6, Corollary 7.6]).
A first interesting application of Theorem 1.4 is the following. Since $\operatorname{Inn}({\check{\Gamma}}^{\pm}(S))$ identifies with a subgroup of $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}(S))$ and the absolute Galois group $G_{\mathbb{Q}}$ embeds in $\operatorname{Out}^{\mathbb{I}}({\check{\Gamma}}(S))$, we have (cf. [27, Proposition 4, (ii)], for a similar result for the profinite Grothendieck-Teichmüller group $\widehat{\operatorname{GT}}$):
Corollary 1.6.
@@
The analogy between the second item of Theorem 1.4 and the corresponding isomorphisms in the topological case $\operatorname{Inn}({\Gamma}^{\pm}(S))=\operatorname{Aut}({\Gamma}^{\pm}(S))\cong\operatorname{Aut}(\mathrm{P}\Gamma^{\pm}(S))$, for $d(S)>1$ and $S\neq S_{1,2},S_{2}$, falls short only in that we do not know whether there holds $\operatorname{Aut}^{\mathbb{I}}({\check{\Gamma}}^{\pm}(S))=\operatorname{Aut}({\check{\Gamma}}^{\pm}(S))$. However, thanks to a recent result by Hoshi, Minamide and Mochizuki (cf. [18]), we are able to fill this gap in genus 0. Since, in this case, we also know that the congruence subgroup property holds (i.e. ${\widehat{\Gamma}}^{\pm}(S)={\check{\Gamma}}^{\pm}(S)$), we get: